A line passing through (–6, –5) and (–1, y) is perpendicular to a line with slope -5/13 . Find the value of y.
Answer by checkley71(8403)

THE SLOPE OF A LINE PERPENDICULAR TO A LINE WITH A SLOPE OF -5/13 IS 13/5.
THEREFORE THE SLOPE OF THE LINE THROUGH (-6,-5) & (-1,Y) IS:
(Y+5)/(-1+6)=13/5
(Y+5)/5=13/5 NOW CROSS MULTIPLY.
5(Y+5)=5*13
Y+5=13
Y=13-5
Y=8 ANSWER.
PROOF
(8+5)/5=13/5
13/5=13/5